### v1

# First Comparative Exoplanetology Within a Transiting Multi-planet System: Comparing the atmospheres of V1298 Tau b and c

11email: s.barat@uva.nl 22institutetext: School of Earth and Planetary Sciences (SEPS), National Institute of Science Education and Research (NISER), Jatani, India 33institutetext: Astrophysics Reseach Center (ARCO), The Open University of Israel, Ra’anana, 43107, Israel 44institutetext: Frontier Research Institute for Interdisciplinary Sciences, Tohoku University, 6-3 Aramaki aza Aoba, Aoba-ku, Sendai, Miyagi 980-8578, Japan 55institutetext: Department of Geophysics, Graduate School of Science, Tohoku University, 6-3 Aramaki aza Aoba, Aoba-ku, Sendai, Miyagi 980-8578, Japan 66institutetext: Institute of Space and Astronautical Science, Japan Aerospace Exploration Agency, 3-1-1 Yoshinodai, Chuo-ku, Sagamihara, Kanagawa 252-5210, Japan 77institutetext: Cluster for Pioneering Research, RIKEN, 2-1 Hirosawa, Wako, Saitama 351-0198, Japan 88institutetext: Department of Astronomy & Astrophysics, University of California, Santa Cruz, CA 95064, USA. 99institutetext: Department of Astronomy and Astrophysics, University of Chicago, Chicago, IL, USA 1010institutetext: School of Earth and Space Exploration, Arizona State University, Tempe, AZ 85287, USA 1111institutetext: Department of Physics, University of Warwick, Coventry CV4 7AL, UK 1212institutetext: Centre for Exoplanets and Habitability, University of Warwick, Coventry CV4 7AL, UK. 1313institutetext: Lowell Observatory, 1400 W Mars Hill Road, Flagstaff, AZ, 86001, USA 1414institutetext: Center for Computational Astrophysics, Flatiron Institute, New York, NY 10010, USA. 1515institutetext: Astrobiology Center, 2-21-1 Osawa, Mitaka, Tokyo 181-8588, Japan. 1616institutetext: National Astronomical Observatory of Japan,2-21-1 Osawa, Mitaka, Tokyo 181-8588, Japan. 1717institutetext: Department of Astronomy, Graduate University for Advanced Studies (SOKENDAI), Tokyo, 2-21-1 Osawa, Mitaka, Japan.
First Comparative Exoplanetology Within a Transiting Multi-planet System: Comparing the atmospheres of V1298 Tau b and c
The V1298 Tau system (20-30Myr), is a benchmark young multi-planet system that provides the opportunity to perform comparative exoplanetology between planets orbiting the same star right after their formation. We present the first atmospheric comparison between two planets in the same transiting system: V1298 Tau b and V1298 Tau c. We derive constraints on the mass of planet b and c (¡20M⊕ at 3 confidence level and respectively) and atmospheric metallicity (logZ/Z⊙=-2.04, -0.16 respectively) from atmospheric retrievals. The V1298 Tau planets, are likely to be similar in terms of mass at the current age, implying that both planets are potential sub-Neptune/super-Earth progenitors. However, planet c is expected to lose a higher fraction of its mass compared to planet b given its close proximity to the host star. Alternatively, the observed spectrum of planet c can be explained by atmospheric hazes, which is in contrast to planet b where efficient haze formation can be ruled out. Higher haze formation efficiency in planet c could be due to differences in atmospheric composition, temperature and higher UV flux incident compared to planet b.
1 Introduction
Demographic studies on mature exoplanets (Fulton et al., 2017; Van Eylen et al., 2018; Szabó & Kiss, 2011) suggest that early evolutionary processes like photoevaporation (Owen & Wu, 2017) and core-powered mass loss (Ginzburg et al., 2018; Gupta & Schlichting, 2019) significantly influence the structure and composition of super-Earths and sub-Neptunes. Therefore, observing planets in their early stages is key to understanding these processes.
V1298 Tau, a 20-30 Myr old weak-lined T-Tauri star in the foreground of the Taurus-Auriga star forming region (Oh et al., 2017; Luhman, 2018), is particularly valuable for such studies. It hosts four transiting planets, making it one of the youngest known multi-planet systems. The inner three planets (V1298 Tau c, d, and b) are in a near 3:2:1 mean-motion resonance (Feinstein et al., 2022; Sikora et al., 2023; David et al., 2019b), suggesting possible migration within a gas disk (Terquem & Papaloizou, 2018; Kajtazi et al., 2023). The orbital period of the outermost planet, V1298 Tau e, remains unconfirmed but is under investigation (Damasso et al., 2023).
In this paper, we compare the atmospheric composition and chemistry of two planets in the V1298 Tau system: V1298 Tau b (0.9 RJ, 24.14 day period) and V1298 Tau c (0.5 RJ, 8.24 day period). A comparison between the known physical properties of both planets is shown in Table 1. V1298 Tau c receives four times the stellar irradiation as V1298 Tau b due to their orbital distance ratio of 1:2. Stellar XUV flux is believed to play an important role in atmospheric chemistry (Kawashima & Ikoma, 2018; Tsai et al., 2023; Venot & Agúndez, 2015; Shulyak et al., 2020) as well as early evolution (Owen & Wu, 2013, 2017), which we aim to test with two planets in the same young system.
Mass measurement for these young planets is challenging due to stellar activity affecting radial velocity (RV) observations (Blunt et al., 2023). Initial RV estimates (Suárez Mascareño et al., 2021), which reported Jovian masses for the two outer V1298 Tau planets, has been challenged (Sikora et al., 2023; Finociety et al., 2023). Barat et al. (2024) estimated V1298 Tau b’s mass to be less than 23M⊕ at 3 level of confidence from HST/WFC3 transmission spectra, and ruled out earlier RV measurements at 5 level of confidence. Ongoing Transit timing variations (TTVs) studies (private communication) suggest that both V1298 Tau b and c are in the sub-Neptune/super-Earth mass regime, indicating they are likely sub-Neptune/super-Earth progenitors with similar masses.
We have observed a transmission spectrum of V1298 Tau c using the same instrumental configuration as Barat et al. (2024) to allow for comparison. We discuss the data reduction, light curve analysis and atmospheric modelling in Section 2, followed by the main results in Section 3. We discuss and interpret our findings in Section 4 and 5.
2 Data analysis and modelling
2.1 Observations
We observed one primary transit of V1298 Tau c (18th October 2021) with 8 HST orbits using the WFC3 instrument for Program GO 16462. The observations were taken with the G141 grism in bidirectional spatial scanning mode covering a range of 1.1-1.7 m. We used the pixel subarray and SPARS25, NSAMP=5 readout mode which resulted in 88.4 s exposures. We used a 0.23”/second scan rate resulting a 170 pixel long image in the spatial direction.
2.2 Data reduction and light curve analysis
We use a custom data reduction pipeline for analyzing the raw HST spatially scanned images (Arcangeli et al., 2018; Jacobs et al., 2022; Barat et al., 2024) mirroring the approach for planet b in Barat et al. (2024).
The broadband integrated light curve is shown in Figure 4. It shows ‘hook’ like systematics, which are typical for similar HST observations (Wakeford et al., 2016). Our observations covered only one pre-transit orbit and no ingress/egress observations. Typically, the first orbit of a visit for HST is known to exhibit large systematics, and it is common practice to not include it in time series analysis (e.g see Kreidberg et al., 2014; Jacobs et al., 2022; Barat et al., 2024). However, since we have only one pre-transit orbit, we do not remove it in our light curve analysis. Further details of the light curve fitting methods are provided in Appendix B.
2.3 Atmospheric modelling with ATMO and planet mass
We interpret the transmission spectra of V1298 Tau c and V1298 Tau b using a chemically consistent ATMO retrieval framework (Tremblin et al., 2015; Drummond et al., 2016; Goyal et al., 2018). ATMO has been previously used to analyze various exoplanet atmospheres (see for e.g Evans et al., 2018; Goyal et al., 2019; Carter et al., 2020). Details about the ATMO models are provided in Appendix D.
Figure 1 shows the best-fit atmospheric models for both planets alongside the observed spectra. The retrieved parameters are listed in Table 2, with posterior distributions in Figures 12, 13. Following Barat et al. (2024), we include planet mass as a free parameter. For planet b, the posterior distribution peaks around 8M⊕ with a 3 upper limit of 20M, consistent with Barat et al. (2024). For planet c, the posterior peaks around 17M⊕ with a 3 upper limit of 50M⊕, consistent with previous RV upper limits (Finociety et al., 2023; Sikora et al., 2023). These results suggest both planets are likely in the sub-Neptune/super-Earth mass range.
Our retrievals indicate low atmospheric metallicity for both planets, with planet b having 0.1-0.01 solar and planet c consistent with solar. The temperature of planet b peaks around 850-900K (See Appendix E for further discussion on planet b’s temperature), while planet c’s temperature is unconstrained. No constraints were found on the C/O ratio or clouds for either planet.
We also performed retrievals with fixed masses to assess their impact on other parameters. For V1298 Tau c, we used 10 and 50M, and for planet b, we used 10 and 23M. The posterior distributions for these fixed mass retrievals are shown in Figures 15, 14, with retrieved parameters in Table 2. The overall conclusion of sub-solar/solar atmospheric metallicity for both planets remains valid even with fixed masses.
3 Results
3.1 The transmission spectrum of V1298 Tau c
We compared the observed transmission spectrum of V1298 Tau c to a flat line and found that models with atmospheric water absorption are slightly favored (2.5). To confirm the presence of water, we conducted a free atmospheric retrieval using the open-source code PetitRadtrans (Mollière et al., 2019). We fixed the planet mass and temperature to the median of the ATMO retrieval posteriors (Table 2) and included a grey cloud deck opacity. The posterior distribution from this fit is shown in Figure 11. While we did not find constraints on CH4, our models preferred the presence of water, with a lower limit for the water volume mixing ratio () at 3 confidence level.
Comparing the spectrum of V1298 Tau c to its sibling planet, V1298 Tau b, revealed that the transit depth uncertainties for V1298 Tau c are larger by a factor of about 1.5. This higher uncertainty is likely due to including the first orbit of the visit for V1298 Tau c, known to have larger wavelength-dependent systematics. Consequently, we did not apply the common-mode correction to the spectroscopic light curves of planet c, as we did for planet b. The common-mode correction could correct for the curvature of the time-series baseline, which then allowed us to fit a linear baseline to the spectroscopic light curves of planet b.
3.2 Relative transmission spectroscopy between planet b and c
In Figure 1 (lower panel), we present the first relative transmission spectrum between two planets within the same system. This relative spectrum is independent of the star’s radius and limb darkening. Stellar inhomogeneities are known to contaminate transmission spectra (Désert et al., 2011; Rackham et al., 2019), which has also been highlighted in recent JWST observations(May et al., 2023; Moran et al., 2023; Lim et al., 2023). A relative transmission spectrum between planets theoretically bypasses stellar contamination issues, assuming consistent stellar contamination across both transits. This enables a better relative comparison of abundances between the planets without precise information on the stellar properties, including metallicity, elemental abundance ratios, and limb-darkening. However, in the case of V1298 Tau, an active star, changes in the stellar surface between observation epochs may still introduce contamination to the relative spectrum extracted at different times.
| Planet | Parameter | Free mass retrieval (ATMO) | Fixed mass (10M⊕) | Fixed mass (50M⊕) |
| logZ [solar] | -0.16 | -0.02 | -0.001 | |
| Rp [] | 0.48 | 0.48 | 0.48 | |
| V1298 Tau c | Teq [K] | 947 | 921 | 985 |
| log | 0.26 | 0.73 | -0.11 | |
| C/O | 0.06 | 0.07 | 0.06 | |
| Mass [M⊕] | 17 | 10 (fixed) | 50 (fixed) | |
| Planet | Parameter | Free mass retrieval (ATMO) | Fixed mass (10M⊕) | Fixed mass (23M⊕) |
| logZ [solar] | -2.04 | -1.80 | -0.86 | |
| Rp [] | 0.84 | 0.84 | 0.84 | |
| V1298 Tau b | Teq [K] | 889 | 806 | 824 |
| log | 1.03 | 0.88 | 0.25 | |
| C/O | 0.08 | 0.05 | 0.04 | |
| Mass [M⊕] | 7.58 | 10 (fixed) | 23 (fixed) |
3.3 Haze formation in the atmosphere of V1298 Tau planets
Hazes have been found to be ubiquitous in exoplanet atmospheres (Kempton et al., 2023; Libby-Roberts et al., 2020a; Knutson et al., 2014). The temperature of V1298 Tau b and c is expected to be between 600-1000K and the host star is young and active, producing copious amounts of XUV radiation, which are important ingredients for haze formation (Yu et al., 2021).
We simulated the transmission spectrum of V1298 Tau b and c, including hazes.We assumed tholin like composition for the hazes (Khare et al., 1984). The simulations were done following the approach in Kawashima & Ikoma (2019). The details of the haze microphysics model are provided in Kawashima & Ikoma (2018). We assumed 20M⊕ and 23M⊕ (Barat et al., 2024) for planet c and b respectively. For further details see Appendix F. The simulated transmission spectrum for V1298 Tau b and c, assuming a 1% haze formation efficiency is shown in Figure 1. The effect of hazes in the HST bandpass is to damp the water absorption feature. Given the large observable water feature in V1298 Tau b’s spectrum, efficient haze formation can be ruled out for this planet. However, atmospheric models with hazes can explain the observed transmission spectrum of planet c. If we assume a lower mass for planet c (10M⊕), which is consistent with current estimates from TTV analysis, we would require even higher haze formation rates to explain the observed spectrum of V1298 Tau c.
4 Discussion
4.1 Relative orbital distances and stellar density
In multi-planet systems, it is possible to fix one planet’s semi-major axis based on another’s measured semi-major axis, since they are related to their orbital periods ratio. However, due to known TTVs in this system, we refrain from fixing the semi-major axis in this analysis. The semi-major axis ratio from independent fits to the white light curves of V1298 Tau b and c is 2.080.1 and agrees with the expected value from their known orbital periods. Assuming circular orbits, the semi-major axis can be used to estimate stellar density (Sandford & Kipping, 2017). From the fits to our white light curves, we estimate the stellar density from the planet b and planet c to be 0.700.01gm/cm3 and 0.670.1gm/cm3. These values align with the published stellar density for V1298 Tau (0.680.1gm/cm3, David et al. (2019b)). indicating that it is inflation compared to similar-mass main sequence stars.
4.2 Limb darkening of V1298 Tau
Modeling stellar limb darkening is crucial for deriving transmission spectra. Prior studies often employed fixed limb darkening for main-sequence stars based on theoretical stellar structure models (e.g, see Deming et al., 2013; Sing et al., 2016). Barat et al. (2024) fitted a linear limb darkening coefficient for V1298 Tau, revealing significant differences compared to theoretical models for a main-sequence star with a similar spectral type (K1). Higher limb darkening compared to theoretical expectations for mature stars has previously been noted for -Pictoris (Landman et al., 2024). The fitted limb darkening coefficients and a theoretical linear limb darkening model is shown in Figure 9. This constitutes the first direct measurement of limb darkening on a weak-lined T-Tauri star.
In Section 4.1, we highlighted that V1298 Tau’s density is lower than that of similar-mass main-sequence stars. Furthermore, it has a relatively high magnetic field strength of the order of 100-300 Gauss (Finociety et al., 2023). High magnetic fields have been proposed as a potential cause for discrepancies between observed and theoretical limb darkening models (Kostogryz et al., 2024). Therefore, a combination of structural differences and higher magnetic fields compared to main-sequence stars, could explain the differences we find with the model.
While leveraging light curves from multiple planets in the same system could potentially reduce limb darkening coefficient uncertainties, the absence of ingress or egress coverage in the light curves of planet c led us to fix the limb darkening coefficients to the values from Barat et al. (2024). A comparison of the transmission spectrum of planet c derived using fixed and free limb darkening is shown in Figure 8. The transit depth uncertainties are higher by 5-10% for each spectroscopic bin in the latter case. Residuals in both cases are approximately 1.2 times the expected photon noise, indicating that stellar variability and instrumental systematics from the first orbit are likely to be the dominant sources of noise in our case.
4.3 Comparative exoplanetology between V1298 Tau b and c
A comparison between the atmospheres of V1298 Tau b and c, measured relative to atmospheric scale height (Figure 1, upper panel), shows that these two planets could host similar atmospheres. Our retrievals show that these planets are likely to have similar masses in the sub-Neptune/super-Earth regime (Table 2). Furthermore, we also find that the atmospheric metallicities of both the planets are solar/sub-solar, which is relatively low compared to their mature counterparts which have been found to have metal-rich/hazy (¿solar) atmospheres (Désert et al., 2011; Kempton et al., 2023; Madhusudhan et al., 2023; Libby-Roberts et al., 2020b). In this scenario, both the V1298 Tau planets are similar in terms of mass and composition, but their observable atmospheres are distinct from their mature counterparts.
Atmospheric hazes can be ruled out for V1298 Tau b, given the large atmospheric absorption feature, however for planet c atmospheric hazes can explain the observed spectrum.. A comparison between the transmission spectra of planet b and c shows that planet c could have a higher haze formation. Haze formation has been linked to the UV flux of the host star; a higher UV flux results in higher monomer production rates for organic hazes (Kawashima & Ikoma, 2019). Higher atmospheric metallicity could also lead to higher haze formation. The higher haze formation in planet c compared to its sibling planet could be due to difference in metallicity, temperature and incident UV flux of the two planets.
4.4 Comparing the early evolution of V1298 Tau b and c
We ran evolutionary models for the V1298 Tau planets. These models start from the time of disk dispersal, and model the structural and thermal evolution of the planet, including atmospheric mass loss due to photoevaporation (Owen & Wu, 2013). We run models assuming both a core-envelope interior structure, as well as with a gradually mixed interior. We have used the evolutionary models presented in Vazan et al. (2022). A brief description of the models is provided in Appendix G. Figure 2 shows a possible mass and radius evolution tracks for these planets.
V1298 Tau c being closer to its host star compared to planet b, is more susceptible to atmospheric mass loss driven by photoevaporation. While the final mass and radius depend on photoevaporation efficiency and the star’s activity level (Poppenhaeger et al., 2021), our models show that V1298 Tau c is likely to lose a larger fraction of its current envelope compared to planet b. Therefore, even if both planets have similar masses now, their early evolution could lead them to different mass-radius regimes.
Our models assume solar metallicity for the envelope opacity. However, the transmission spectrum of V1298 Tau c could also be explained using atmospheric hazes, unlike V1298 Tau b (Figure 1. Young planets are expected to cool and contract over time (Kubyshkina et al., 2020; Linder et al., 2019). The presence of high altitude hazes can increasing the atmospheric opacity, therefore increasing the cooling and contraction timescale (Lee & Chiang, 2015), leading to higher photoevaporative mass loss. Thus, a hazy atmosphere of V1298 Tau c as opposed to a clear atmosphere of V1298 Tau b, could lead to slower cooling and higher mass loss.
4.5 Comparing the composition of V1298 Tau planet with the exoplanet population
Figure 3 shows V1298 Tau b and c on the mass-metallicity and metallicity-orbital distance planes. The observed low atmospheric metallicity for these planets deviates from the mass-metallicity trend predicted by core-accretion models (Pollack et al., 1996; Bodenheimer & Pollack, 1986) and reported for exoplanets (Welbanks et al., 2019; Thorngren & Fortney, 2019). This could potentially be explained by ongoing evolutionary processes that transform their current metal-poor atmospheres into the metal-rich atmospheres seen in mature sub-Neptune/super-Earth planets (Désert et al., 2011; Kempton et al., 2023; Libby-Roberts et al., 2020a; Madhusudhan et al., 2023).
We compare the V1298 Tau system with the 30 Myr old HR 8799 system (Marois et al., 2010) with four directly imaged planets (HR8799 b, c , d and e) with masses 7, 10, 10, 10 (Marois et al., 2010; GRAVITY Collaboration et al., 2019). Nasedkin et al. (2024) using ground based spectra from the VLTI/GRAVITY program, report super-stellar(100 times stellar) atmospheric metallicity for these planets. Assuming a -0.65 dex metallicity for HR8799 (Swastik et al., 2021) we convert the retrieved atmospheric metallicity to stellar unit.
The HR 8799 planets exhibit higher atmospheric metallicity compared to the mass-metallicity trend (Thorngren & Fortney, 2019) predicted by core-accretion formation, while V1298 Tau planets fall below this trend. HR 8799 is a system of widely separated massive gas giants, whereas V1298 Tau likely consists of compact, low-mass sub-Neptune/super-Earth progenitors. Both systems likely host primordial atmospheres. The different mass regimes suggest HR 8799 planets underwent runaway gas accretion or gravitational collapse, unlike V1298 Tau planets. Theories propose close-in sub-Neptunes and super-Earths formed in situ within the water ice-line in dusty or depleting disks to prevent runaway gas accretion (Lee et al., 2014; Lee & Chiang, 2016). Wide orbit gas giants likely formed beyond the ice lines, allowing for runaway accretion (Pollack et al., 1996) or gravitational instability (Boss, 1997; Dodson-Robinson et al., 2009), but these mechanisms alone do not explain HR 8799’s high metallicity, necessitating the accretion of volatile-rich solids (Nasedkin et al., 2024). Formation location relative to ice lines could lead to differences in volatile enrichment, as beyond the ice-line, the disk contains volatile-rich solids, enhancing the metal content of primordial atmospheres (Öberg et al., 2011; Bitsch et al., 2019).
4.6 Formation and evolution of multi-planet systems
The mass of the V1298 Tau planets indicates they are likely sub-Neptune/super-Earth progenitors, which are in a near mean-motion resonance (1:2:3). Atmospheric retrievals suggest these planets could have low metallicity atmospheres. Population synthesis studies (Mordasini, 2018; Benz et al., 2014) show sub-Neptunes and super-Earths often form in multi-planet systems in mean motion resonance (Lambrechts et al., 2019). Emsenhuber et al. (2023) classified systems with multiple low-mass planets into two types: Class I: inner dry/rocky planets formed in situ with outer icy sub-Neptunes migrating inwards, and Class II: multiple sub-Neptunes forming beyond the water ice line and migrating inwards. In the first scenario, inner planets have volatile-poor atmospheres, whereas, in the second, all planets are volatile-rich.
The relatively low atmospheric metallicity of V1298 Tau b and c suggests a formation similar to class I planets . However, these young planets’ atmospheric composition may evolve. In the pebble accretion scenario, a metallicity gradient forms in the protoplanet’s envelope (Ormel et al., 2021; Bodenheimer et al., 2018), with metal-rich layers hidden deep inside. Over time, atmospheric mass loss can strip volatile-poor upper layers, enriching atmospheric metallicity (Vazan et al., 2022), aligning them with mature sub-Neptune/super-Earth population. In this case, the V1298 Tau system could also have a formation similar to the class II systems.
If V1298 Tau c has a hazy atmosphere, it might be more metal-rich than the outer planet b, which doesn’t fit either class I or class II systems. Class I systems have inner volatile-poor planets, while class II systems have planets with similar compositions. Due to its proximity to the star, V1298 Tau c is likely to lose more of its primordial envelope compared to V1298 Tau b 2. Thus, V1298 Tau b and c might have formed like class II systems, with volatiles hidden deep inside, but atmospheric mass loss caused planet c’s composition to evolve faster than planet b, explaining their difference in their atmospheric composition.
5 Conclusion
In this paper, we perform the first comparative study of atmospheres of multiple planets in the same transiting system. Comparing the transmission spectra of the young V1298 Tau b and c planets reveals that planet b has a large, haze-free envelope, while the presence of hazes in planet c cannot be ruled out. Atmospheric models suggest similar masses for both planets in the sub-Neptune/super-Earth regime. These models also find relatively low atmospheric metallicities for both these young planets compared to their matured counterparts, indicating they are sub-Neptune/super-Earth progenitors. Alternatively, if planet c has a hazy atmosphere, higher haze formation in planet c compared to planet b could be due to differences in composition, temperature, and incoming stellar UV flux. The higher envelope opacity of planet c would affect its cooling rate and evolution differently than for planet b.
6 Acknowledgements
J.M.D acknowledges support from the Amsterdam Academic Alliance (AAA) Program, and the European Research Council (ERC) European Union’s Horizon 2020 research and innovation program (grant agreement no. 679633; Exo-Atmos). This work is part of the research program VIDI New Frontiers in Exoplanetary Climatology with project number 614.001.601, which is (partly) financed by the Dutch Research Council (NWO). Y.K. acknowledges the support from JSPS KAKENHI Grant Numbers JP21K13984 and 22H05150.
7 Softwares used
numpy, scipy, matplotlib, emcee, lmfit, pandas, ATMO, PetitRadtrans
Appendix A Description of raw data reduction
We start our analysis from the ‘ima.fits’ files downloaded from MAST archive. During each exposure, the detector is read multiple times, without flushing out the accumulated charge (non-destructive reads). We create sub-exposures by subtracting consecutive non-destructive reads for each exposure which are reduced separately. This technique has been widely applied in past observations with HST (e.g, see Kreidberg et al. 2014; Arcangeli et al. 2018; Jacobs et al. 2022). We apply a wavelength dependent flat-field correction and flag bad pixels with data quality DQ=4, 32, or 512 by calwf3. We apply a local median filter to identify cosmic rays and clip pixels which are more than five median deviations from the median. To account for the dispersion direction drift of the spectrum (Wakeford et al. 2016) we use the first exposure of a visit as a template and shift the spectrum for an exposure along the dispersion direction to match the template. The maximum shift that we require is 0.1 pixels. We apply optimal extraction algorithm (Horne 1986) on each subexposure to maximize signal-to-noise ratio.
Appendix B Light curve fitting
HST WFC3 light curves often exhibit strong systematics like charge-trapping ramps and visit-long slopes (Kreidberg et al. 2014; Deming et al. 2013; Tsiaras et al. 2016; Wakeford et al. 2016). Common practice involves discarding the first orbit due to significantly larger systematics (See Figure 4) compared to the rest of the visit (e.g see, Wakeford et al. 2013; Arcangeli et al. 2018; Jacobs et al. 2022; Barat et al. 2024). However, in this case, we could not remove from the analysis because the transit occurred two hours earlier than expected from the linear ephemeris (Feinstein et al. 2022) due to large TTVs and we could only secure one orbit prior to ingress.
The first-orbit systematics are strongly wavelength dependent (Zhou et al. 2020, 2017) and cannot be corrected for using common-mode approaches, which had been previously applied to similar observations of V1298 tau b (Barat et al. 2024). Therefore, to handle these systematics, we used the physically motivated RECTE charge-trapping model (Zhou et al. 2017), which fits the first orbit ramp better than other methods (Berta et al. 2012; Kreidberg et al. 2014). We fixed the mid-transit time using simultaneous TESS observations (Feinstein et al. 2022). V1298 Tau shows strong rotational variability, with an average amplitude of 2 and a rotational period of 2.83 days (David et al. 2019b). We modeled it with a second-order polynomial for the out-of-transit baseline. See Appendix C for more discussion on the baseline modelling. The planetary transit signal was modelled using batman (Kreidberg 2015).
For the white light curve (1.12-1.6 m), we fit the RECTE parameters and the baseline polynomial coefficients, along with the planet’s transit depth and semi-major axis. Other orbital parameters were fixed to literature values (Feinstein et al. 2022; Sikora et al. 2023; David et al. 2019a). The RECTE parameters constitute 4 parameters: the number of initially filled fast/slow charge traps and the number of fast/slow charge traps that are filled in between orbits. Due to the lack of ingress and egress data, our limb darkening precision was impaired. However, by using the limb darkening parameter retrieved from the observations of planet b (Barat et al. 2024), we were able to significantly reduce the uncertainties on the transit spectrum (Figure 8). This highlights the advantage and uniqueness of studying multiple planets in the same system. For a discussion on the limb darkening of V1298 Tau see Appendix 4.2. We also independently measure the stellar density from the HST white light curves which is discussed in Appendix 4.1.
The fit was performed using emcee (Foreman-Mackey et al. 2013). The posterior distribution from these fits is shown in Supplementary Figure 5 and the best-fit transit model and residuals, are shown in Supplementary Figure 4 (upper panel).
Spectroscopic light curves were analyzed similarly by dividing the stellar spectra into 7-pixel bins, resulting in 17 channels (identical to V1298 Tau b). We fixed the semi-major axis from the white light curve fits and fit the RECTE parameters and baseline for each bin, and the transit depth. We fix the wavelength dependent limb-darkening coefficients to values derived in Barat et al. (2024). The detrended spectroscopic light curves and residuals are shown in Supplementary Figure 6.
Appendix C Modelling the baseline
Young stars exhibit quasi-periodic flux modulation due to rotational variability (e.g see Feinstein et al. 2022; David et al. 2019b; Mann et al. 2016, 2022). The transit of planet c analyzed in this work occurred near an inflection point of the stellar baseline, resulting in significant deviation from a linear function. We modeled the baseline using linear, quadratic, and cubic polynomial functions for all the spectroscopic channels. The transit depths from all three fits were consistent within 1, but quadratic and cubic polynomials were statistically favored over a linear baseline (See Supplementary Figure 7).
Higher-order polynomials, however, introduce correlation between the transit depth and baseline, increasing uncertainties and scatter in the transmission spectrum. The Bayesian Information Criterion (BIC) did not improve significantly with a cubic baseline compared to a quadratic one, and the cubic polynomial coefficients strongly correlated with the transit depths, likely adding noise. Therefore, we used a quadratic polynomial to model the baseline for the final transmission spectrum.
Appendix D Atmospheric models with ATMO
ATMO’s python based retrieval framework was applied and bench-marked in Lewis et al. (2020) and Rathcke et al. (2021) for HAT-P-41b and WASP-79b, respectively . This same retrieval framework is applied to V1298 Tau b and c in this work. This framework includes a Nested sampling Bayesian sampler DYNESTY (Speagle 2020) coupled to ATMO. We perform chemical equilibrium retrieval with isothermal - profile. Chemical equilibrium abundances are computed using Gibbs energy minimization as detailed in Drummond et al. (2016). We include H2-H2, H2-He Collision Induced opacity (CIA), as well as H2O, CH4 and CO high temperature molecular opacity in the retrieval. The clouds are treated as grey cloud deck between 1 and 50 millibar pressure levels, but with variable cloud opacity governed by the cloud factor used as the variable in the retrieval, as detailed in Goyal et al. (2018).
Appendix E Comparing V1298 Tau b ATMO retrieval with Barat et al. (2024)
For this study, we performed atmospheric retrieval on V1298 Tau b using the same setup as for planet c, assuming an isothermal atmosphere. Our results, indicating sub-solar metallicity and a mass upper limit of 20M⊕, align with Barat et al. (2024), who used PetitRadtrans. However, our retrieved temperatures are higher (800-900K) compared to their 600-700K. The discrepancy arises because Barat et al. (2024) used a non-isothermal temperature-pressure (T-P) profile with high interior temperature and vertical mixing to explain the non-detection of methane for this planet. Our isothermal model lacks this flexibility, leading to higher temperatures beyond the CH4/CO transition (Fortney et al. 2020) to eliminate methane from the upper atmosphere.
Appendix F Haze formation models
The haze microphysics models were simulated using the models described in Kawashima & Ikoma (2018), following the approach of Kawashima & Ikoma (2019). First, using a photochemical model the abundance of gaseous species was calculated. The XUV spectrum of V1298 Tau was adopted from Duvvuri et al. (2023). Haze production rates (or efficiency of haze formation) are uncertain, therefore we assumed a certain fraction of precursor molecules to be dissociated to form haze monomers. The haze precursor molecules are assumed to be CH4, HCN and C2H2 in our model. We derived vertical density and radius distributions of the haze particles following Kawashima & Ikoma (2018). We assumed solar elemental abundance ratio and other parameters such as eddy diffusion coefficient, monomer radius, internal density, and refractive index of haze particles similar to Kreidberg et al. (2022). We simulated transmission spectra, for V1298 Tau b and c assuming 23 and 20M⊕ respectively with obtained profiles of gas and haze particles. In Figure 1 we show the simulated transmission spectra assuming a 1 haze production efficiency.
Appendix G Theoretical evolutionary models
Our models start from the time of disk dispersal and calculate the thermal evolution and structural evolution. We consider both core-envelope as well as gradually mixed interior structures. We include atmospheric mas loss due to photoevaporation from the topmost layer of the planets using the formalism of Rogers & Owen (2021). The mass loss rates are estimated from an energy-limited model (Erkaev et al. 2007), assuming a fixed photoevaporation efficiency. Our models include core erosion and mixing with the envelope. The evolutionary models are described in Vazan et al. (2022). We choose the initial conditions (mass, radius) such that it matches the current mass and radius for the planets. We assumed 23M⊕ and 20M⊕ for the current mass estimates of the V1298 Tau b and c respectively. We run simulations with both core-envelope and gradually mixed interior models. Both interior structure models can explain the current mass/radius of the planets. Mass and radius (normalized at current value) evolution models for the V1298 Tau planets are shown in Figure 2.
Appendix H Effect of stellar activity on transmission spectrum
We estimate the effect of stellar activity on the transmission spectrum of V1298 Tau c. We adopt the method outlined in Rackham et al. (2019), where we estimate a correction factor based on the spot coverage fraction and spot temperature contrast. We adopt a spot coverage fraction of 20% for V1298 Tau (Feinstein et al. 2021). We adopt an extreme spot contrast of 1000K following spot contrasts derived for a sample of T Tauri stars (Koen 2016). We show a comparison between the observed and corrected transmission spectrum in Supplementary Figure 10. The observed and contamination corrected normalized spectra are consistent within 1 for all spectroscopic channels.
Since V1298 Tau is a K-type star, its contamination function does not show molecular features like M-dwarfs (Barclay et al. 2021). Therefore, the effect of the stellar correction is to change the level of the continuum in this case. However, the atmospheric features, which are relative to the continuum remain largely unaffected.
Appendix I Supplementary Figures
| Planet name | Mass [M⊕] | Atmospheric metallicity [solar] | Semi-major axis [AU] | Stellar metallicity [dex] | Reference |
| Wasp-18b | 10.200.35 | 1.03 | 0.02 | 0.110.08 | Coulombe et al. (2023) |
| Wasp-39b | 0.2800.03 | 10 | 0.048 | -0.030.1 | Rustamkulov et al. (2023) |
| Wasp-77Ab | 1.660.06 | 0.12 | 0.023 | -0.010.1 | August et al. (2023) |
| Wasp-96b | 0.480.03 | 7.5 | 0.045 | 0.140.19 | Radica et al. (2023) |
| TOI-270d | 0.0150.001 | 225 | 0.07 | -0.20.1 | Benneke et al. (2024) |
| Wasp-107b | 0.10.005 | 10 | 0.05 | 0.020.09 | Welbanks et al. (2024) |
| HD149026b | 0.3580.01 | 20 | 0.0436 | 0.330.1 | Gagnebin et al. (2024) |
| HD209458b | 0.59029040.01 | 3 | 0.05 | -0.010.05 | Xue et al. (2024) |
| Wasp-80b | 0.530.03 | 6 | 0.03 | -0.10.1 | Bell et al. (2023) |
| Wasp-17b | 0.780.2 | 50 | 0.051 | -0.070.1 | Grant et al. (2023) |
| GJ1214b | 0.0260.001 | 100 | 0.01 | 0.020.1 | Kempton et al. (2023) |